### v3

@@ 3.2 Multi-domain reconstruction
The primary advantage of reconstructing $\vec{M}_{rec}$ is to allow the visualisation of local variations in magnetic domains and the twisting of magnetic domain walls. Fig. 6a and Fig. 3c are projected along the $z$ axis. The close-to-anti-parallel $\vec{M}_{rec}$ vectors in Fig. 6a explain the phase shift variations at the nanowire intersection in Fig. 3c. The $\text{curl}(\vec{M}_{rec})$ direction in Fig. 6b shows that in Fig. 3c only part of the vortex has its curl oriented close to the $z$ axis, while the other part is more in-plane and requires imaging about a second tilt axis for a sufficient signal to be detected. The magnetic vortex core is curved to favour its positioning at the protrusion at the bottom of the intersection, which is considered to be a domain wall pinning site. The limited spatial resolution is evident in Fig. 6c, where the vortex core has a lower value of $M_{s}$ than the surrounding material due to spatial averaging of vector components of opposite sign. This averaging is a result of the regularisation in Eq. 5, in which the preferential selection of solutions with low gradients introduces smoothing. Although this smoothing is necessary to constrain solutions for $\vec{M}_{rec}$ to retain ferromagnetic ordering, the averaging of anti-parallel vectors offsets $M_{s}$ values at the vortex core. In the present experiment, the $M_{s}$ offset due to spatial averaging is negligible when compared to variations in material composition. The elemental map shown in Fig. 3d indicates that the cobalt content at the intersection is elevated to $70\pm 15$%, with a predicted $\mu_{0}M_{s}~{}=~{}0.8\pm 0.5$ T (Teresa et al., 2016; Pablo-Navarro et al., 2018; Donnelly et al., 2022), in close agreement with the value of $\mu_{0}M_{s}~{}=~{}0.9\pm 0.3$ T measured from the reconstruction. The magnetic domain wall occupies the full $\sim 200$ nm width of the intersection, which is comparable to the $\sim 100$ nm magnetic domain wall width of transverse head-to-head domain walls that previously observed using magnetic X-ray laminography in 80% purity FEBID cobalt nanowire double helixes (Donnelly et al., 2022). In summary, MBIR has provided accurate measurements of $M_{s}$ for a U-shaped vortex magnetic domain wall of width 200 nm, but is less sensitive to sub-50 nm $M_{s}$ variations in the present study due to limitations in spatial resolution and sensitivity that can be overcome in future experiments.
@@ 4.2 Outlook
Previous works have shown 2D reconstructions of $\vec{M}$ (Kovács et al., 2017; Song et al., 2018), while an alternative MBIR implementation (Mohan et al., 2018) has provided only a qualitative 3D reconstruction. Micromagnetic simulations (Wolf et al., 2022; Lyu et al., 2024) are a common approach to 3D quantification, but require a knowledge of the magnetic properties of the material, whereas MBIR can, in special cases, reconstruct $\vec{M}$ without such prior knowledge. Quantitative measurements have been reported using both X-ray and neutron tomography (Donnelly et al., 2017; Hilger et al., 2018; Donnelly et al., 2020, 2022; Di Pietro Martínez et al., 2023), with a spatial resolution of 50 nm for X-ray laminography. When applied with improved spatial resolution, MBIR would be uniquely suited to characterise 3D magnetic nanostructures, such as nanowire junctions, interfaces in nanoparticles, Bloch-type skyrmions on curved surfaces, and hopfions. In total, this method reconstructs projections of $\vec{B}$, projections of electrostatic potential, 3D distributions of $\vec{M}$, and 3D geometry of nanostructures.
@@ 5 Conclusions
Model-based iterative reconstruction has been applied and further developed to measure the 3D distribution of magnetisation $\vec{M}$ in a FEBID ferromagnetic nanostructure from multiple projections of the magnetic induction field recorded using off-axis electron holography. The effect of null space limitations on the uniqueness and accuracy of the reconstruction has been assessed. Optimal estimation diagnostics show that the reconstruction has a precision of 0.01 T at each voxel. The present spatial resolution of several tens of nanometres is restricted primarily by computational limitations and can be improved greatly in future experiments (towards the atomic scale) by using a smaller voxel size, more specialised electron microscopes, and improved image processing.
@@ 6 Availability of data and materials
Data is available on request. Software and example datasets are available from repositories listed in the Supplementary Material [S3].

### v4

@@ 3.2 Multi-domain reconstruction
The primary advantage of reconstructing $\vec{M}_{rec}$ is to allow the visualisation of local variations in magnetic domains and the twisting of magnetic domain walls. Fig. 6a and Fig. 3c are projected along the $z$ axis. The close-to-anti-parallel $\vec{M}_{rec}$ vectors in Fig. 6a explain the phase shift variations at the nanowire intersection in Fig. 3c. The $\text{curl}(\vec{M}_{rec})$ direction in Fig. 6b shows that in Fig. 3c only part of the vortex has its curl oriented close to the $z$ axis, while the other part is more in-plane and requires imaging about a second tilt axis for a sufficient signal to be detected. The magnetic vortex core is curved to favour its positioning at the protrusion at the bottom of the intersection, which is considered to be a domain wall pinning site. The limited spatial resolution is evident in Fig. 6c, where the vortex core has a lower value of $M_{s}$ than the surrounding material due to spatial averaging of vector components of opposite sign. This averaging is a result of the regularisation in Eq. 5, in which the preferential selection of solutions with low gradients introduces smoothing. Although this smoothing is necessary to constrain solutions for $\vec{M}_{rec}$ to retain ferromagnetic ordering, the averaging of anti-parallel vectors offsets $M_{s}$ values at the vortex core. In the present experiment, the $M_{s}$ offset due to spatial averaging is negligible when compared to variations in material composition. The elemental map shown in Fig. 3d indicates that the cobalt content at the intersection is elevated to $70\pm 15$%, with a predicted $\mu_{0}M_{s}~{}=~{}0.8\pm 0.5$ T (Teresa et al., 2016; Pablo-Navarro et al., 2018; Donnelly et al., 2022), in close agreement with the value of $\mu_{0}M_{s}~{}=~{}0.9\pm 0.3$ T measured from the reconstruction. The correlation between $M_{s}$ and the atomic Co content is consistent with previous work, as shown in Supplementary Material [S6]. The magnetic domain wall occupies the full box200similar-toabsent200$\sim$ 200∼ 200 nm width of the intersection, which is comparable to the $\sim 100$ nm magnetic domain wall width of transverse head-to-head domain walls that previously observed using magnetic X-ray laminography in 80% purity FEBID cobalt nanowire double helixes (Donnelly et al., 2022). In summary, MBIR has provided accurate measurements of $M_{s}$ for a U-shaped vortex magnetic domain wall of width 200 nm, but is less sensitive to sub-50 nm $M_{s}$ variations in the present study due to limitations in spatial resolution and sensitivity that can be overcome in future experiments.
@@ 4.2 Outlook
Previous works have shown 2D reconstructions of $\vec{M}$ (Kovács et al., 2017; Song et al., 2018), while an alternative MBIR implementation (Aditya Mohan et al., 2018) has provided only a qualitative 3D reconstruction. Micromagnetic simulations (Wolf et al., 2022; Lyu et al., 2024) are a common approach to 3D quantification, but require a knowledge of the magnetic properties of the material, whereas MBIR can, in special cases, reconstruct $\vec{M}$ without such prior knowledge. Quantitative measurements have been reported using both X-ray and neutron tomography (Donnelly et al., 2017; Hilger et al., 2018; Donnelly et al., 2020, 2022; Di Pietro Martínez et al., 2023), with a spatial resolution of 50 nm for X-ray laminography. Our MBIR results are consistent with X-ray tomography results and micromagnetic simulations of similar FEBID nanostructures (Donnelly et al., 2022). When experimental and computational limitations are resolved to improve spatial resolution, MBIR would be uniquely suited to characterise 3D magnetic nanostructures, such as nanowire junctions, interfaces in nanoparticles, Bloch-type skyrmions on curved surfaces, and hopfions. In total, this method reconstructs projections of $\vec{B}$, projections of electrostatic potential, 3D distributions of $\vec{M}$, and 3D geometry of nanostructures.
@@ 5 Conclusions
Model-based iterative reconstruction has been applied and further developed to measure the 3D distribution of magnetisation $\vec{M}$ in a FEBID ferromagnetic nanostructure from multiple projections of the magnetic induction field recorded using off-axis electron holography. The effect of null space limitations on the uniqueness and accuracy of the reconstruction has been assessed. The results are in agreement with X-ray tomography results obtained from similar structures. Optimal estimation diagnostics show that the reconstruction has a precision of 0.01 T at each voxel. The present spatial resolution of several tens of nanometres is restricted primarily by computational limitations and can be improved greatly in future experiments (towards the atomic scale) by using a smaller voxel size, more specialised electron microscopes, and improved image processing.
@@ 6 Availability of data and materials
The data underlying this article are available in Enlighten at https://doi.org/10.5525/gla.researchdata.1953.